OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + 2y = 0 2x + 4y = 0 one solution no solution coincident

OpenStudy (anonymous):

@ribhu

OpenStudy (amorfide):

if you look closely, you see that if you multiply equation 1 by 2 that it becomes the same line as equation 2 therefore you have infinite solutions

OpenStudy (anonymous):

Solve the linear system. x + y = -3 y = 2x (-1, -2) no solution coincident

OpenStudy (ribhu):

sorry coincident

OpenStudy (anonymous):

What about this one?

OpenStudy (ribhu):

coincident

OpenStudy (anonymous):

Both answers are coincident?

OpenStudy (ribhu):

last one one solution, this one coincident

OpenStudy (anonymous):

Thought you said no solution?

OpenStudy (ribhu):

did i? wait

OpenStudy (ribhu):

one solution for first and coincident for this one

OpenStudy (anonymous):

x + y = 6 x - y = 8 Solve the system of equations. (1, -7) (7, -1) no solution

OpenStudy (anonymous):

???

OpenStudy (anonymous):

The first one has infinite solutions, because it comes out to be 0 = 0 The second one has one solution, (-1, -2) Solved by substitution. Your third one is substitution again, one solution, (7,-1)

OpenStudy (anonymous):

Please ask any questions if you're not sure how I got those answers.

OpenStudy (anonymous):

@Kommander_Kitten But infinite solutions isn't one of the answer choices?